# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

### v2

@@ 1 Introduction
Conditions ($A_{1}$)($A_{2}$) ensure that the equations admit the symmetries of the product group $O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, while conditions ($A_{3}$) and ($A_{4}$) imply the well-behavedness of our non-linearity at solutions to (1) near zero and infinity, respectively. Specifically, ($A_{2}$) allows the equivariant degree to distinguish between stationary and non-stationary functions, ($A_{3}$) guarantees that the nonlinearity does not interact with the spectrum of $A$ and the Nagumo condition ($A_{4}$) is essential for establishing a-priori bounds on the magnitude of solutions to (1). System (1) is a natural model for the dynamics of a configuration of $N$ identical pendula, subjected to nonlinear forcing $f$ and with coupling relations specified by the matrix $A$ and the group $\Gamma$. In this context, the vector $x(t):=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t))$ represents the angles of the pendula from the vertical position, $\dot{x}(t):=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t))$ the angular velocities and $\ddot{x}(t):=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t))$ the angular accelerations at any time $t>0$. Methods based on the Leray-Schauder degree have been used to solve a wide variety of second order differential equations [2, 9, 14, 17, 18, 22, 24]. However, the inability of the Leray-Schauder degree to distinguish between periodic solutions that differ by a fixed time shift and differentiate constant solutions from non-constant solutions have hindered its suitibility for autonomous differential equations. The effectiveness of the $G$-equivariant Leray-Schauder degree in detecting non-constant solutions to the system (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [13]. More generally, the $G$-equivariant Leray-Schauder degree has been applied to problems admitting appropriate spatial-temporal symmetries [3, 4, 11, 12, 15, 19]. Compared with other techniques to solve systems similar to (1), the equivariant degree approach requires relatively minimal regularity assumptions on the nonlinearity. For example, existence results were obtained in [1] by topological index argument and in [10] using Morse theory framework, under the assumption that the right-hand-side interacts with only finitely many eigenvalues of the differential operator. A prerequisite for the application of any $G$-equivariant degree-theory based argument to (1) is the reformulation of our system as a fixed point equation in an appropriate functional $G$-space with a nonlinear operator in the form of a $G$-equivariant compact perturbation of identity. In [13], Garcia-Azpetia et. al prove that the problem of finding $p$-periodic solutions to (1) is equivalent to the problem of finding zeros of the operator
@@
where $\beta:=\frac{p}{2\pi}$, $\mathscr{H}:=H^{2}_{2\pi}([0,2\pi];V)$ and $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{H}$ is the differential operator $\mathscr{L}u:=\ddot{u}-u$. Under the conditions ($A_{1}$)($A_{4}$) and provided that the linearization $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, Garcia-Azpetia et. al additionally demonstrate, that:
@@ 3. $\rm(iii)$
The exact number of irreducible $\Gamma$-representations $\mathcal{V}_{j}^{-}$ ‘contained’ in the $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is called the $\mathcal{V}_{j}$-isotypic multiplicity of $V$ and is calculated according to the ratio,
@@
$e^{i\theta}w:=e^{im\theta}w,\quad\kappa w:=\bar{w},\quad\theta\in SO(2),\;\kappa\in O(2),\;w\in\mathcal{W}_{m},$
and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially such that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},\quad\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$ (3)
@@
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. Under the assumptions ($A_{0}$)($A_{4}$), if the matrix $A:V\rightarrow V$ admits an odd number of eigenvalues $\mu_{j}\in\sigma(A)$ satisfying
@@ 2 Setting for the Leray-Schauder $G$-Equivariant Degree
Notice that, with the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the system (1) become:
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi),\end{cases}$ (4)
such that any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (4) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original equation. Following [13], we denote by $\mathscr{H}$ the Sobolev space of $2\pi$-periodic, $V$-valued functions
@@ 2.1 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
On the other hand, each degree $G\text{\rm-deg}(\mathscr{A}_{m,j}(\alpha),B(\mathscr{H}_{m,j}))$ is fully specified by the $\mathcal{V}_{j}$-isotypic multiplicities $\{m_{j}\}_{j=0}^{r}$ together with the real spectra of $\mathscr{A}_{m,j}$ according to formula,
@@
which which contribute non-trivially to $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$. To begin, the negative spectrum of $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ is accounted for with the index set
@@ 3 Orbit Types of Maximal Kind
$\operatorname{coeff}^{H}((K)\cdot(L))=\frac{n(H,L)|W(L)|n(H,K)|W(K)|-\sum_{(\tilde{H})>(H)}\operatorname{coeff}^{\tilde{H}}((K)\cdot(L))n(H,\tilde{H})|W(\tilde{H})|}{|W(H)|},$
is only useful if one first identifies the set of all orbit types $(\tilde{H})\in\Phi_{0}(G)\setminus\{(G)\}$ with $(\tilde{H})>(H)$. Therefore, computation of the coefficients $\operatorname{coeff}^{H}((K)\cdot(L))$ is generally impractical except for those orbit types which are maximal in $\Phi_{0}(G)$, i.e. for those orbit types $(H)\in\Phi_{0}(G)$ where $(\tilde{H})>(H)$ implies $(\tilde{H})=(G)$, in which case the recurrence formula simplifies to
@@
Motivated by the Remark 3.1, the conjugacy class of a subgroup $H\leq G$ is called an orbit type of maximal kind if it is a maximal element in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some $m>0$. Moreover, we denote by $\mathfrak{M}$ the set of all orbit types of maximal kind, by $\mathfrak{M}_{m}$ the set of maximal elements in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\mathfrak{M}_{m}\cap\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$. Since each orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ is contained in the sublattice of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{0,1,\ldots,r\}$, we have
@@ 3.1 The $s$-Folding Homomorphism
Given any orbit type $(H)\in\Phi_{0}(G)$ and a number $s_{0}\in{\mathbb{N}}$, we will demonstrate how the map (16) can be used to to describe the set
$\mathfrak{N}(H,s_{0}):=\{s\in{\mathbb{N}}:({}^{s}H)\leq({}^{s_{0}}H)\}.$ (17)
@@
$\mathfrak{m}(H)\mid s_{0}-s_{1}\text{ or }\mathfrak{m}(H)\mid s_{0}+s_{1}.$
@@
With these notations, it becomes clear that the subgroups ${}^{s_{0}}H,{}^{s_{1}}H\leq G$ have the relation (19) for some $g=(a,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ if and only if $\varphi^{s_{0}}_{O}(axa^{-1})=\varphi^{s_{1}}_{O}(x)$ for every $(x,y)\in K_{O}\times K_{\Gamma}$ with $\varphi_{O}(x)=\varphi_{\Gamma}(y)$ such that the relation (18) is equivalent to the existence of an element $a\in O(2)$ satisfying
$\varphi^{s_{1}}_{O}(axa^{-1})=\varphi^{s_{0}}_{O}(x)\text{ for all }x\in K_{O}.$
Clearly, it is sufficient to consider the behaviour of the rotation generator in $r\in K_{0}$ and the two cases $a\in\{\kappa,e_{O(2)}\}$. If $a=\kappa$, then from $\kappa r\kappa=r^{-1}$, one has
$\kappa{}^{s_{1}}H\kappa\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(\kappa r\kappa)=\varphi_{O}^{s_{0}}(r)\iff\varphi_{O}(r)^{s_{0}+s_{1}}=e_{L}\iff\mathfrak{m}(H)\mid s_{0}+s_{1}.$
@@
${}^{s_{1}}H\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(r)=\varphi_{O}^{s_{0}}(r)\iff\varphi_{O}(r)^{s_{0}-s_{1}}=e_{L}\iff\mathfrak{m}(H)\mid s_{0}-s_{1}.$
@@
$[P]:=\begin{cases}1\quad&\text{ if }P\text{ is true};\\ 0\quad&\text{ otherwise}.\end{cases}$ (20)
@@
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=\begin{cases}|W(H)|\quad&\text{ if }s=\gcd(s_{0},s_{1})\text{ and }s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1});\\ 0\quad&\text{otherwise},\end{cases}$
@@
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=|W(H)|[s=\gcd{(s_{0},s_{1})}][s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})].$
@@
${}^{s_{0}}H\cap g{}^{s_{1}}Hg^{-1}$ =$=\left\{(x,y)\in\left({}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}\right)\times K_{\Gamma}:\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(a^{-1}xa)=\varphi_{\Gamma}(y)\right\}$
@@
The equivalence of condition $(\rm i)$ with the inclusion $s\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$ has already been demonstrated in the proof of Theorem 4.2. On the other hand, the observation
@@
implies that condition $(\rm ii)$ can be equated with $s=\gcd(s_{0},s_{1})$ (cf. [6], Example 3.12). For convenience, put $d:=\gcd(s_{0},s_{1})$ and notice that, in the case one has $d\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$, the coefficient $n_{d}\in{\mathbb{Z}}$ standing next to the orbit type $({}^{d}H)$ in the Burnside Ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G)$ can be obtained with the recurrence formula (cf. Appendix A) as follows
@@
and the result follows from the fact $|W(H)|=|W({}^{s}H)|$ for any $s\in{\mathbb{N}}$ together with the observation that $d\in\mathfrak{N}(H,s_{0})\cap\mathfrak{N}(H,s_{1})$ implies $n({}^{d}H,{}^{s_{0}}H)=n({}^{d}H,{}^{s_{1}}H)=1$. $\square$
@@
Since, the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside Ring products of the form (11). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the $G$-equivariant Leray-Schauder degree (cf. Appendix A) implies
@@
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{1}}^{H}\text{ and }\dim\mathcal{V}_{m,j_{2}}^{H}\text{ are of the same parity;}\\ -x_{0}\quad&\text{else,}\end{cases}$
@@
where in either of the cases: $x_{0}=2$ and $|W(H)|=1$ or $x_{0}=1$ and $|W(H)|=2$, one has $02-x_{0}|W(H)|=$ 0. Suppose instead that $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are of different parties and one has the two corresponding cases, $y_{m,j_{1}}=x_{0}$ and $y_{m,j_{2}}=0$ or $y_{m,j_{1}}=0$ and $y_{m,j_{2}}=x_{0}$, both of which imply $y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}$. $\square$ Since, our computations in the Burnside Ring often involve the product of a number of basic degrees, we consider a natural generalization of Lemma (5.2) to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and the Burnside Ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$.
@@
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{k}}^{H}\text{ is odd for an even number of }j_{k}\in\{0,1,\ldots,r\};\\ -x_{0}\quad&\text{otherwise,}\end{cases}$
@@
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and $s_{1},s_{2}\in{\mathbb{N}}$ where $s_{1}\neq s_{2}$ and for any number any $s_{0}\in{\mathbb{N}}$, one has
@@
$\quad+2x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}\dim\mathcal{V}_{m,j_{2}}^{H}\right]\left[\operatorname{gcd}(s_{1},s_{2})=s_{0}\right][s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)],$

### v3

@@ 1 Introduction
Conditions ($A_{1}$)($A_{2}$) ensure that the equations admit the symmetries of the product group $O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, while conditions ($A_{3}$) and ($A_{4}$) imply the well-behavedness of our non-linearity at solutions to (1) near zero and infinity, respectively. Specifically, ($A_{2}$) allows the equivariant degree to distinguish between stationary and non-stationary functions, ($A_{3}$) guarantees that the nonlinearity does not interact with the spectrum of $A$ and the Nagumo condition ($A_{4}$) is essential for establishing a-priori bounds on the magnitude of solutions to (1). System (1) is a natural model for the dynamics of a configuration of $N$ identical pendula, subjected to nonlinear forcing $f$ and with coupling relations specified by the matrix $A$ and the group $\Gamma$. In this context, the vector $x(t):=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t))$ represents the angles of the pendula from the vertical position, $\dot{x}(t):=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t))$ the angular velocities and $\ddot{x}(t):=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t))$ the angular accelerations at any time $t>0$. Methods based on the Leray-Schauder degree have been used to solve a wide variety of second order differential equations [3, 9, 14, 17, 18, 22, 24]. However, the inability of the Leray-Schauder degree to distinguish between periodic solutions that differ by a fixed time shift and differentiate constant solutions from non-constant solutions have hindered its suitability for autonomous differential equations. The effectiveness of the $G$-equivariant Leray-Schauder degree in detecting non-constant solutions to the system (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [13]. More generally, the $G$-equivariant Leray-Schauder degree has been applied to problems admitting appropriate spatial-temporal symmetries [4, 5, 11, 12, 15, 19]. Compared with other techniques to solve systems similar to (1), the equivariant degree approach requires relatively minimal regularity assumptions on the nonlinearity. For example, existence results were obtained in [2] by topological index argument and in [10] using Morse theory framework, under the assumption that the right-hand-side interacts with only finitely many eigenvalues of the differential operator. A prerequisite for the application of any $G$-equivariant degree-theory based argument to (1) is the reformulation of our system as a fixed point equation in an appropriate functional $G$-space with a nonlinear operator in the form of a $G$-equivariant compact perturbation of identity. In [13], Garcia-Azpetia et al. prove that the problem of finding $p$-periodic solutions to (1) is equivalent to the problem of finding zeros of the operator
@@
where $\beta:=\frac{p}{2\pi}$, $\mathscr{H}:=H^{2}_{2\pi}([0,2\pi];V)$ and $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{H}$ is the differential operator $\mathscr{L}u:=\ddot{u}-u$. Under the conditions ($A_{1}$)($A_{4}$) and provided that the linearization $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, Garcia-Azpetia et al. additionally demonstrate that:
@@ 3. $\rm(iii)$
The exact number of irreducible $\Gamma$-representations $\mathcal{V}_{j}^{-}$ ‘contained’ in the $\Gamma\times\mathbb{Z}_{2}$-isotypic component $V_{j}$ is called the $\mathcal{V}_{j}$-isotypic multiplicity of $V$ and is calculated according to the ratio:
@@
$e^{i\theta}w:=e^{im\theta}w,\quad\kappa w:=\bar{w},\quad e^{i\theta},\kappa\in O(2),\;w\in\mathcal{W}_{m},$
(here, $\kappa$ indicates the reflection generator in $O(2)$) and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially such that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},$ (3)
where
$\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$
@@
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. Under the assumptions ($A_{0}$)($A_{4}$), if the matrix $A:V\rightarrow V$ admits an odd number of eigenvalues $\mu_{j}\in\sigma(A)$ with odd isotypic multiplicity $m_{j}$ (cf. (2)) satisfying
@@ 2 Setting for the Leray-Schauder $G$-Equivariant Degree
Notice that, with the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the system (1) becomes
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi).\end{cases}$ (4)
Clearly, any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (4) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original equation. Following [13], we denote by $\mathscr{H}$ the Sobolev space of $2\pi$-periodic, $V$-valued functions
@@ 2.1 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
On the other hand, each degree $G\text{\rm-deg}(\mathscr{A}_{m,j},B(\mathscr{H}_{m,j}))$ is fully specified by the $\mathcal{V}_{j}$-isotypic multiplicities $\{m_{j}\}_{j=0}^{r}$ together with the real spectra of $\mathscr{A}_{m,j}$ according to formula:
@@
which contribute non-trivially to $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$. To begin, the negative spectrum of $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ is accounted for with the index set
@@ 3 Orbit Types of Maximal Kind
$C^{H}(K,L)=$
$\quad\frac{n(H,L)|W(L)|n(H,K)|W(K)|-\sum_{(\tilde{H})>(H)}C^{\tilde{H}}(K,L)n(H,\tilde{H})|W(\tilde{H})|}{|W(H)|},$
(here, we have temporarily adopted the notation $C^{H}(K,L):=\operatorname{coeff}^{H}((K)\cdot(L))$ due to constraints on space) is only useful if one first identifies the set of all orbit types $(\tilde{H})\in\Phi_{0}(G)\setminus\{(G)\}$ with $(\tilde{H})>(H)$. Therefore, computation of the coefficients $\operatorname{coeff}^{H}((K)\cdot(L))$ is generally impractical except for those orbit types which are maximal in $\Phi_{0}(G)$, i.e. for those orbit types $(H)\in\Phi_{0}(G)$ where $(\tilde{H})>(H)$ implies $(\tilde{H})=(G)$, in which case the recurrence formula simplifies to
@@
Motivated by Remark 3.1, the conjugacy class of a subgroup $H\leq G$ is called an orbit type of maximal kind if it is a maximal element in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some $m>0$. Moreover, we denote by $\mathfrak{M}$ the set of all orbit types of maximal kind, by $\mathfrak{M}_{m}$ the set of maximal elements in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\mathfrak{M}_{m}\cap\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$. Since each orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ is contained in the sublattice of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{0,1,\ldots,r\}$, we have
@@ 3.1 The $s$-Folding Homomorphism
Given any orbit type $(H)\in\Phi_{0}(G)$ and a number $s_{0}\in{\mathbb{N}}$, we will demonstrate how the map (16) can be used to describe the set
$\left\{s\in{\mathbb{N}}:({}^{s}H)\leq({}^{s_{0}}H)\right\}.$ (17)
@@
$\mathfrak{m}(H)\mid\frac{s_{0}-s_{1}}{\gcd(s_{0},s_{1})}\text{ or}\mathfrak{m}(H)\mid\frac{s_{0}+s_{1}}{\gcd(s_{0},s_{1})}.$
@@
With these notations, it becomes clear that the subgroups ${}^{s_{0}}H,{}^{s_{1}}H\leq G$ have the relation (19) for some $g=(a,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ if and only if $\varphi^{s_{0}}_{O}(axa^{-1})=\varphi^{s_{1}}_{O}(x)$ for every $(x,y)\in{}^{\gcd(s_{0},s_{1})}K_{O}\times K_{\Gamma}$ satisfying $\varphi^{\gcd(s_{0},s_{1})}_{O}(x)=\varphi_{\Gamma}(y)$. In particular, since the $\Gamma$ component of an amalgamated subgroup is unchanged by folding, the relation (18) is equivalent to the existence of an element $a\in O(2)$ satisfying
$\varphi^{s_{1}}_{O}(axa^{-1})=\varphi^{s_{0}}_{O}(x)\text{ for all }x\in{}^{d}K_{O},\text{ where }d:=\gcd(s_{0},s_{1}).$
In turn, it is sufficient to only consider the behavior of the element $r^{\frac{1}{d}}\in{}^{d}K_{O}$ where $r$ is the rotation generator in $K_{O}$ and the two cases $a\in\{\kappa,e_{O(2)}\}$. If $a=\kappa$, then from $\kappa r\kappa=r^{-1}$, one has
$\kappa{}^{s_{1}}H\kappa\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(\kappa r^{\frac{1}{d}}\kappa)=\varphi_{O}^{s_{0}}(r^{\frac{1}{d}})\iff\varphi_{O}(r)^{\frac{s_{0}+s_{1}}{d}}=e_{L}\iff\mathfrak{m}(H)\mid\frac{s_{0}+s_{1}}{d}.$
@@
${}^{s_{1}}H\leq{}^{s_{0}}H\iff\varphi_{O}^{s_{1}}(r^{\frac{1}{d}})=\varphi_{O}^{s_{0}}(r^{\frac{1}{d}})\iff\varphi_{O}(r)^{\frac{s_{0}-s_{1}}{d}}=e_{L}\iff\mathfrak{m}(H)\mid\frac{s_{0}-s_{1}}{d}.$
@@
$[P]:=\begin{cases}1\quad&\text{ if }P\text{ is true};\\ 0\quad&\text{ otherwise},\end{cases}$ (20)
and for any set of indices $I\subset{\mathbb{N}}$, we indicate by $\mathcal{B}_{\mathfrak{m}}(I)$ the Boolean expression
[…]
$\mathfrak{m}|\frac{(x-y)}{\gcd(x,y)}\text{ or }\mathfrak{m}|\frac{(x+y)}{\gcd(x,y)}.\text{"}$ ”
@@
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=\begin{cases}|W(H)|\quad&\text{ if }s=\gcd(s_{0},s_{1})\text{ and }[\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1}\})];\\ 0\quad&\text{otherwise},\end{cases}$
@@
$\operatorname{coeff}^{{}^{s}H}(({}^{s_{0}}H)\cdot({}^{s_{1}}H))=|W(H)|[s=\gcd{(s_{0},s_{1})}][\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1}\})].$
@@
${}^{s_{0}}H\cap$ $g\,{}^{s_{1}}Hg^{-1}$
=$=\left\{(x,y)\in\left({}^{s_{0}}K_{O}\cap a{}^{s_{1}}K_{O}a^{-1}\right)\times K_{\Gamma}:\varphi_{O}^{s_{0}}(x)=\varphi_{O}^{s_{1}}(a^{-1}xa)=\varphi_{\Gamma}(y)\right\}$
@@
The equivalence of condition $(\rm i)$ with the truthfulness of the Boolean expression $\mathcal{B}_{\mathfrak{m}(H)}(\{s,s_{0},s_{1}\})$ has already been demonstrated in the proof of Theorem 4.2. On the other hand, the observation
@@
implies that condition $(\rm ii)$ can be equated with $s=\gcd(s_{0},s_{1})$ (cf. [7], Example 3.12). For convenience, put $d:=\gcd(s_{0},s_{1})$ and notice that, in the case one has $\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{0},s_{1}\})$, the coefficient $n_{d}\in{\mathbb{Z}}$ standing next to the orbit type $({}^{d}H)$ in the Burnside ring product $({}^{s_{0}}H)\cdot({}^{s_{1}}H)\in A(G)$ can be obtained with the recurrence formula (cf. Appendix A) as follows
@@
and the result follows from the fact $|W(H)|=|W({}^{s}H)|$ for any $s\in{\mathbb{N}}$ together with the observation that $\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{0},s_{1}\})$ implies $n({}^{d}H,{}^{s_{0}}H)=n({}^{d}H,{}^{s_{1}}H)=1$. $\square$
@@
Since the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside ring products of the form (11). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the $G$-equivariant Leray-Schauder degree (cf. Appendix A) implies
@@
$\operatorname{coeff}^{H}(\deg_{\mathcal{V}_{m,j_{1}}}\cdot\deg_{\mathcal{V}_{m,j_{2}}})=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{1}}^{H}\text{ and }\dim\mathcal{V}_{m,j_{2}}^{H}\\ &\quad\text{ are of the same parity;}\\ -x_{0}\quad&\text{else,}\end{cases}$
@@
where in either of the cases: $x_{0}=2$ and $|W(H)|=1$ or $x_{0}=1$ and $|W(H)|=2$, one has $02-x_{0}|W(H)|=$ 0. Suppose instead that $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are of different parties and one has the two corresponding cases, $y_{m,j_{1}}=x_{0}$ and $y_{m,j_{2}}=0$ or $y_{m,j_{1}}=0$ and $y_{m,j_{2}}=x_{0}$, both of which imply $y_{m,j_{1}}+y_{m,j_{2}}-y_{m,j_{1}}y_{m,j_{2}}|W(H)|=x_{0}$. $\square$ Since our computations in the Burnside ring often involve the product of a number of basic degrees, we consider a natural generalization of Lemma (5.2) to an orbit type of maximal kind $(H)\in\mathfrak{M}$ and the Burnside ring product of any finite collection of basic degrees $\{\deg_{\mathcal{V}_{m,j_{k}}}\}_{k=1}^{N}$ with $m>0$ and $j_{1},\ldots,j_{N}\in\{0,1,\ldots,r\}$.
@@
$\operatorname{coeff}^{H}\left(\prod\limits_{k=1}^{N}\deg_{\mathcal{V}_{m,j_{k}}}\right)=\begin{cases}0\quad&\text{if }\dim\mathcal{V}_{m,j_{k}}^{H}\text{ is odd }\\ &\quad\text{for an even number of }j_{k}\in\{0,1,\ldots,r\};\\ -x_{0}\quad&\text{otherwise,}\end{cases}$
@@
Let $(H)\in\mathfrak{M}$ be an orbit type of maximal kind. For any pair of basic degrees $\deg_{\mathcal{V}_{s_{1}m,j_{1}}},\deg_{\mathcal{V}_{s_{2}m,j_{2}}}\in A(G)$ with $m>0$, $j_{1},j_{2}\in\{0,1,\ldots,r\}$ and $s_{1},s_{2}\in{\mathbb{N}}$ where $s_{1}\neq s_{2}$ and for any number $s_{0}\in{\mathbb{N}}$, one has
@@
$\quad+2x_{0}\left[2\nmid\dim\mathcal{V}_{m,j_{1}}^{H}\dim\mathcal{V}_{m,j_{2}}^{H}\right]\left[\operatorname{gcd}(s_{1},s_{2})=s_{0}\right][\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})],$